kingraid:

jed rode his scooter to school at 20 mph and then jagged back at 8mph if the round trip took him 7 hours how far was it to school?

SmokeyBrown:

Hi, and welcome to QuestionCove! We're trying to solve for distance, and we have a few values given to us. The total time of the trip is 7 hours. This includes the time it took Jed to scooter the distance added to the time it took to jog the distance. The time taken for each part of the trip is equal to the distance divided by the speed Jed was going. Using this information, we can set up the equation: 7 hours = distance / (20mph) + distance (8mph) or simplifying out the units and using variables to make it easier to read: 7 = d/20 + d/8 We can use algebra combine like terms and solve for d: 7 = 2d/40 + 5d/40 7 = 7d/40 280 = 7d 40 = d So, the distance from the starting point and school is 40 miles. We can check our work by checking how long this trip would take: Travelling 40 miles on scooter (20mph) takes 2 hours (40/20=2). Travelling 40 miles while jogging (8mph) takes 5 hours (40/8=5). 2 hours + 5 hours = 7 hours, which is how long it took, so it looks like we got the right answer :)